Opened 4 years ago

Last modified 2 years ago

#503 new flight processing

THL13-01, flight day 267/2013, Thales

Reported by: knpa Owned by: edfi
Priority: immediate Milestone: 2013 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/arsf_internal/GB13_00-2013_267_Thales

Data arrived from ARSF via network transfer on 01/10/2013

GB13/00 Scientific objective: Commercial flight testing the Thales instrument

GB13/00 Priority: urgent

PI: ARSF Internal

Sensors:

Camera (Not requested by PI)
Eagle (Dispatched 30/10/2013)
Hawk (Dispatched 30/10/2013)
LiDAR (Not requested but useful for checks)

Change History (18)

comment:1 Changed 4 years ago by knpa

  • Description modified (diff)
  • Summary changed from GB13/00, flight day 267/2013, Thales to THL13-01, flight day 267/2013, Thales

comment:2 Changed 4 years ago by knpa

  • Description modified (diff)

comment:3 Changed 4 years ago by edfi

Starting Navigation processing

comment:4 Changed 4 years ago by edfi

PPP_base

Sano (Isle of Wight)

Lat 50 39 01.19399
Lon -1 12 46.70872
Hei 91.107

comment:5 Changed 4 years ago by edfi

Processing Hyperspectral

Line 4 missing dark frames, used dark frames from line 5

comment:6 Changed 4 years ago by edfi

Lines 5 & 6 loaded onto ftp ARSF15 (04/10/13)

comment:7 Changed 4 years ago by edfi

Lines 123 PPP (Farb)

51 16 47.77148
-0 46 21.69130
112.689

comment:8 Changed 4 years ago by edfi

This flight "A" is taken from splitting up the original Thales flight. The data was split up according the flight lines. So far this has been done to the hyperspectral files and the RCD. The RCD raw files were kept/deleted according to the frame number from the log sheets. Further more the event files were edited to reflect the images which were kept. Files which were not immediately obvious were to trim were kept in tact. Where files have been edited an original copy should be also located within the folder.

comment:9 Changed 4 years ago by tipo

All LiDAR have zero returns.

comment:10 Changed 4 years ago by stgo

Edit: Wrong ticket.

Last edited 4 years ago by stgo (previous) (diff)

comment:11 Changed 4 years ago by tipo

Created hyperspectral delivery.

Last edited 4 years ago by tipo (previous) (diff)

comment:12 Changed 4 years ago by stgo

Check apl failed looking for e267011b_mask.bil, I have fixed this and remade the readme. Ready to deliver.

comment:13 Changed 4 years ago by tipo

Hyperspectral dispatched 30/10/2013

comment:14 Changed 4 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)

comment:15 Changed 4 years ago by tipo

Created RCD delivery.

comment:16 Changed 4 years ago by stgo

Beginning RCD D/C.

comment:17 Changed 4 years ago by stgo

RCD D/C complete, no issues found.

comment:18 Changed 2 years ago by stgo

Archiving
Started archiving

  • flight didn't need archiving
Last edited 2 years ago by stgo (previous) (diff)
Note: See TracTickets for help on using tickets.